Do you believe that the management of a public company (i.e. listed firm on a stock
exchange) is able to manage the firm’s earnings and overstate its financial
performance through the change of accounting estimates related to the depreciation
policy? Explain your answer and provide examples of a hypothetical and/or a real
firm that engaged in such accounting practices. Please also provide a brief discussion of the reasons that may explain the behavior and the accounting choices of the firm’s
top management in the above described context. (Word Limit 500 words)
Answer #1

I do believe that the Management of a public company is able to manage the firm's earnings and overstate its financial performance. Although, Laws in many countries have changed to provide the companies with the adoption of stipulated useful life, stipulated residual value and further with the method of accounting.

Factors that affect the calculation of rate of depreciation are as follows:

1. Cost of acquisition of asset,

2. Useful life of the asset,

3. Residual value,

4. Method of depreciation used.

Assuming that in the given financial reporting framework entity has an option to choose the method of depreciation to be applied. Change of this single factor affects the amount of depreciation and at the end the financial results.

Take a hypothetical example wherein company listed on a stock exchange has purchased an asset costing $ 100000,

now further consider the residual value is estimated at 10% of cost and life of the asset is expected to be 5 years.

Depreciation under Straight Line Method (SLM) method p.a. would be = 100000 - 10000/5 = 18000 p.a.

whereas depreciation under Written down value method would be calculated 1 - (Residual Value/Cost)^1/n

i.e. 1 -(10000/100000)^1/5 =1-910000/100000

i.e. 1 - 0.6309 = .3691 i.e. 36.91%

Calculating the depreciation & WDV for 5 Years under Two different methods

Year S.L.M. W.D.V
Op.Balance Depreciation Cl. Balance Op.Balance Depreciation Cl.Balance
1 100000 18000 82000 100000 36910 63090
2 82000 18000 64000 63090 23286.52 39803.48
3 64000 18000 46000 39803.48 14691.46 25112.01
4 46000 18000 28000 25112.01 9268.84 15843.17
5 28000 18000 10000 15843.17 5847.71 9995.45

Now it can be seen from the above calculations that deprecation has a significant difference under both the methods given all other conditions are same. In the above situation reporting entity will be benefitted by following SLM method since the depreciation charge will be lower and results will appear to be more appreciable.

If the entity changes its estimate of useful life (given the fact the financial reporting framework that the entity follows allows the entity to estimate useful life of asset rather than prescribing it.) and if the entity estimates useful life to be more then it will lead to lesser depreciation in each year and eventually improve the financial results of the entity.

Similarly, if the entity estimates the residual value (given the fact that the financial reporting framework that the entity follows allows and does not restrict the residual value estimation to any particular amount) to be more again the depreciation charge per annum gets reduced and the entity's financial results will improve.

In the reasons for Top Management's behavior with respect to choosing accounting estimates relating to depreciation.

1. It significantly affects the performance of the entity, wherever entity wants to improve its performance, it will follow accounting estimates which lead to lesser depreciation.

2. Depreciation is the non cash expense, change in depreciation will not affect cashflows but certainly if the depreciation charge is reduced perfomance seems to be improved.